The following problems need to be changed into slope-intercept form:
13x - 11y = -12
9x - 7y = -7
x - 3y = 6
4x - y = 1

Answer by math-vortex(648)

Hi, there--

THE PROBLEM:
The following problems need to be changed into slope-intercept form: 
13x - 11y = -12
9x - 7y = -7
x - 3y = 6
4x - y = 1

A SOLUTION:
All these equations are in standard form. This is the procedure for changing any equation to slope-intetcept form:

(a) Rewrite 13x - 11y = -12 in slope-intercept form.
Isolate the y-term on the left side. (Subtract 13x from both sides.)
13x - 11y = -12
13x  - 11y - 13x = -13x - 12

Combine like terms (13x-13x=0).
-11y = -13x - 12

We want the equation in "y=" form. 
Divide both sides by the coefficient of the y-term. (Divide by -11.)
y = (-13/-11)x - (12/-11)

Simplify the equation.
y = (13/11)x + (12/11)

(b) Rewrite 9x - 7y = -7 in slope-intercept form.
Isolate the y-term on the left hand side. (Subtract 9x from both sides.)
9x - 7y = -7
9x - 7y - 9x = -9x - 7

Combine like terms (9x-9x=0).
- 7y = -9x - 7

We want the equation in "y=" form.
Divide both sides by the coefficient of the y-term. (Divide by -7.)
y = (-9/-7)x - (7/-7)
y = (9/7)x +1

I've worked through the first two. Now you try. Feel free to email if you have any questions about the solution.

Good luck,

Mrs. F
math.in.the.vortex@gmail.com